1. Use the data in the contingency table to answer the question.

Columns
Rows 1 2 3 Total
1 36 35 92 163
2 67 57 113 237
Total 103 92 205 400

You wish to test the null hypothesis of "independence"—that the probability that a response falls in any one row is independent of the column it falls in—and you plan to use a chi-square test. You are given that there are 2 degrees of freedom associated with the test statistic Χ2. You are also given the value of the test statistic Χ2 = 3.16.

Find the rejection region for α = 0.05. (Round your answer to two decimal places.)

Χ2 > ____________

2. Conduct the appropriate test of the specified probabilities using the given information. Use α = 0.05.

The five categories are equally likely to occur, and the category counts are shown in the table

Category 1 2 3 4 5
Observed Count 46 65 73 53

63

Find the test statistic. (Round your answer to two decimal places.)

Χ2=______________
Answer #1

1)

for 2 df and 0.05 level of signifcance critical region       χ2= 5.991

rejection region x2 >5.99

2)

           relative observed Expected residual Chi square
category frequency(p) Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
1 0.200 46.000 60.00 -1.81 3.267
2 0.200 65.000 60.00 0.65 0.417
3 0.200 73.000 60.00 1.68 2.817
4 0.200 53.000 60.00 -0.90 0.817
5 0.200 63.000 60.00 0.39 0.150
total 1.000 300 300 7.4667
test statistic X2 = 7.47